answersLogoWhite

0

The probability of rolling two prime numbers on a standard pair of dice is 1 in 4, or 0.25. Take the probability of rolling a prime on one die, 3 in 6, or 1 in 2, or 0.5, and square it.

User Avatar

Wiki User

∙ 15y ago
User Avatar

What else can I help you with?

Related Questions

What is the probability of not rolling prime numbers on both dice?

It is 1/4.


What is the probability of rolling a prime number on a 6 face dice?

If the numbers are 1 to 6, there are three prime numbers in that range, a probability of 50%.


When rolling two dice what are the odds of rolling both even numbers?

The probability of rolling two even numbers on two standard dice is 0.52, or 0.25.


What is the probability that you get a prime number with one dice and a tail on a coin?

The prime numbers from 1-6 are 2, 3, 5 (1 is not prime). There are 3 prime numbers on a dice and 6 total. therefore the probability of rolling a prime is 3/6. The probability of getting a tails when flipping a coin is 1/2. Therefore you just multiply the two. 3/6 * 1/2 = 1/4


What is the probability of rolling three even numbers on six dice?

The probability is 5/16 = 0.3125


What is the probability of rolling a prime number on a role of dice?

50% 1/2 0.5


What is the probability of rolling a prime number on a six sided dice?

The probability is 1 out of two. The primes you can roll are 2,3, and 5.


Odds Rolling prime number sum with roll of two dice?

The probability is 5/12.


What is the probability of rolling two numbers whose sum is 2 equal?

For 6 sided dice, there is only 1 way to get a 2: (1,1). There are 36 outcomes rolling 2 dice; so the probability of rolling two numbers whose sum is 2 is 1/36.


What is the probability of rolling a 5 on a 6 sided dice?

There is always a 1 in 6 chance of rolling ANY number on a six-sided dice, as there are 6 numbers.


Probability of rolling doubles with two dice?

The probability of rolling doubles with two dice is 1 in 6, or about 0.167.


What are the probability of getting prime number with a dice?

A standard six-sided die has the numbers 1 through 6. The prime numbers within this range are 2, 3, and 5. Therefore, there are 3 prime numbers out of 6 possible outcomes. The probability of rolling a prime number is thus 3/6, which simplifies to 1/2 or 50%.

Trending Questions
Does a steep curve on the line graph indicate a rapid or slow rate of change? What is used as a measure of total risk? How the beta of a portfolio can equal the market beta if 50 percent of the portfolio is invested in a security that has twice the amount of systematic risk as an average risky security? Is 40 percent alcohol a lot? How heavy is th average 11 year old? If you spin 2 spinners once how many possible outcomes are there? If I were to flip a coin 4 times. What is the probability of me getting heads twice? What is the average life span of the 1.00 bill? How much does 250 grams of chicken weigh in imperial measurement? What are the place values of a 14 digit number? What would an example of a null and alternative hypothesis look like for a one right tail test with a rejection region at an alpha of 1 percent? What are some benefits of representing data sets using graphs of frequency distribution? Could you be pregnant if you had your period twice in one month and than have it once the next month? What is the meaning of biggitiy? What is 1760000 rounded to the nearest hundred thousand? What does a bar graph measure? Indicate whether these passage contain a faulty move from correlation to cause.if so state your criticism1.There is a significant correlation between going to the hospital and dyingso hospital are imp? Researchers are studying yield of a crop in two locations The researchers are going to compare two independent 90 percent confidence intervals for the mean yield in each location The probability that? Why might estimation be better than an exact answer? I had my period from Feb 1-5 and then I got it again on Feb 16 is it possible that I'm pregnant?